### v1

@@ 2.4 The clique relaxation
Consider a hypergraph ${\mathcal{G}}=(V,{\mathcal{E}})$ and let $V^{\prime}$ be a subset of $V$. We define the subhypergraph of ${\mathcal{G}}$ induced by $V^{\prime}$ as the hypergraph ${\mathcal{G}}_{V^{\prime}}$ with node set $V^{\prime}$ and with edge set $\{e\cap V^{\prime}:e\in{\mathcal{E}},|e\cap V^{\prime}|\geq 2\}$. For every edge $e$ of ${\mathcal{G}}_{V^{\prime}}$, there may exist several edges $e^{\prime}$ of ${\mathcal{G}}$ satisfying $e=e^{\prime}\cap V$; we denote by $e^{\prime}(e)$ one such arbitrary edge of ${\mathcal{G}}$. For ease of notation, we identify an edge $e$ of ${\mathcal{G}}_{V^{\prime}}$ with an edge $e^{\prime}(e)$ of ${\mathcal{G}}$. Define
$L_{\bar{V}}:=\{z\in{\mathbb{R}}^{V+{\mathcal{E}}}:z_{v}=1\ \forall v\in V\setminus\bar{V}\}.$ (20)
Denote by ${\rm proj}_{{\mathcal{G}}_{V^{\prime}}}(\text{MP}({\mathcal{G}})\cap L_{V^{\prime}})$ the set obtained from $\text{MP}({\mathcal{G}})\cap L_{V^{\prime}}$ by projecting out all variables $z_{v}$, for all $v\in V\setminus V^{\prime}$, and $z_{f}$, for all $f\in{\mathcal{E}}\setminus\{e^{\prime}(e):e\in{\mathcal{E}}_{V^{\prime}}\}$, where ${\mathcal{E}}_{V^{\prime}}$ denotes the edge set of ${\mathcal{G}}_{V^{\prime}}$. In [11], the authors prove the following equivalence:
Lemma 4 (Lemma 6 in [11]).
Let ${\mathcal{G}}=(V,{\mathcal{E}})$ be a hypergraph and let $L_{V^{\prime}}$ be a set defined by (20) for some $V^{\prime}\subseteq V$. Then $\text{MP}({\mathcal{G}}_{V^{\prime}})={\rm proj}_{{\mathcal{G}}_{V^{\prime}}}(\text{MP}({\mathcal{G}})\cap L_{V^{\prime}})$.
The next proposition characterizes UGM hypergraphs for which $\text{MP}^{\rm cl}({\mathcal{G}})=\text{MP}({\mathcal{G}})$.
@@
Let ${\mathcal{G}}=\cup_{C\in{\mathcal{C}}}{{\mathcal{G}}_{C}}$ be a UGM hypergraph where ${\mathcal{C}}$ denotes the set of maximal clique of the binary UGM and ${\mathcal{G}}_{C}$ is a complete hypergraph with node set $C$. Then $\text{MP}^{\rm cl}({\mathcal{G}})=\text{MP}({\mathcal{G}})$ if and only if ${\mathcal{C}}$ has the running intersection property.
@@
Denote by $C_{1},C_{2},\cdots,C_{m}$ a running intersection ordering of ${\mathcal{C}}$. Let ${\mathcal{G}}_{i}$ denote the complete hypergraph with node set $C_{i}$ for all $i\in\{1,\cdots,m\}$. Define ${\mathcal{G}}^{\prime}_{m}=\cup_{i=1}^{m-1}{{\mathcal{G}}_{i}}$. Then by definition of the running intersection ordering, ${\mathcal{G}}_{m}\cap{\mathcal{G}}^{\prime}_{m}$ is a section hypergraph of ${\mathcal{G}}_{i}$ for some $i\in\{1,\cdots,m-1\}$ and hence is a complete hypergraph. Hence, by Theorem 1 the set $S({\mathcal{G}})$ is decomposable into $S({\mathcal{G}}_{m})$ and $S({\mathcal{G}}^{\prime}_{m})$. By a recursive application of this argument $m$ times we conclude that $S({\mathcal{G}})$ is decomposable into multilinear sets $S({\mathcal{G}}_{i})$, $i\in\{1,\cdots,m\}$, implying $\text{MP}^{\rm cl}({\mathcal{G}})=\text{MP}({\mathcal{G}})$.
Now suppose that the set ${\mathcal{C}}$ does not have the running intersection property. This in turn implies that the edge set of ${\mathcal{G}}$ does not have the running intersection property either. Denote by $V$ the node set of ${\mathcal{G}}$. By Lemma 4 to prove $\text{MP}({\mathcal{G}})\subset\text{MP}^{\rm cl}({\mathcal{G}})$, it suffices to show that $\text{MP}({\mathcal{G}}_{V^{\prime}})\subset\text{MP}^{\rm cl}({\mathcal{G}}_{V^{\prime}})$ for some $V^{\prime}\subseteq V$. Let $V^{\prime}\subseteq V$ be such that the edge set ${\mathcal{E}}^{\prime}$ of ${\mathcal{G}}_{V^{\prime}}$ does not have the running intersection property, while the edge set of ${\mathcal{G}}_{V^{\prime\prime}}$ for any $V^{\prime\prime}\subset V^{\prime}$ has the running intersection property. First, note that ${\mathcal{G}}_{V^{\prime}}$ must contain $\beta$-cycles as otherwise ${\mathcal{E}}^{\prime}$ has the running intersection property. Second ${\mathcal{G}}_{V^{\prime}}$ does not contain an edge $e=V^{\prime}$ as otherwise ${\mathcal{E}}^{\prime}$ again has the running intersection property. It then follows that ${\mathcal{G}}_{V^{\prime}}$ is a graph that is a chordless cycle as otherwise there exists $V^{\prime\prime}\subset V^{\prime}$ such that ${\mathcal{G}}_{V^{\prime\prime}}$ a chordless cycle implying that its edge set does not have the running intersection property. The inclusion $\text{MP}({\mathcal{G}}_{V^{\prime}})\subset\text{MP}^{\rm cl}({\mathcal{G}}_{V^{\prime}})$ then follows since odd-cylce inequalities are facet-defining for $\text{MP}({\mathcal{G}}_{V^{\prime}})$ [36] and are not implied by $\text{MP}^{\rm cl}({\mathcal{G}}_{V^{\prime}})$, which only contains McCormick type inequalities. ∎
@@ 2.5 The multi-clique relaxation
By Proposition 3, the clique relaxation is the strongest LP relaxation introduced so far. Yet by Proposition 4, this LP is guaranteed to solve the original problem if the set of cliques ${\mathcal{C}}$ has the running intersection property; a property that is often not present in applications. In this section we propose stronger LP relaxations for Problem B-UGM by constructing the multilinear polytope of a UGM hypergraph containing multiple cliques that do not have the running intersection property. More precisely, we consider a special structure, that we refer to as the cycle of cliques, and obtain the multilinear polytope using disjunctive programming [2]. This structure appears in applications such as image restoration and decoding problems.
Consider the set of maximal cliques ${\mathcal{C}}:=\{C_{1},C_{2},\ldots,C_{m}\}$, where $m\geq 3$ and $|C_{i}|\geq 3$ for all $i\in[m]:=\{1,\cdots,m\}$. We say that ${\mathcal{C}}$ is a cycle of cliques if $C_{i}\cap C_{i+1}=\{v_{i},\bar{v}\}$ for all $i\in[m]$, where $v_{i}\neq v_{j}$ for any $i\neq j\in[m]$ and where we define $C_{m+1}:=C_{1}$. It is simple to check if ${\mathcal{C}}$ is a cycle of cliques, then it does not have the running intersection property. Figure 1 illustrates examples of cycles of cliques that appear in applications. The objective of this section is to characterize $\text{MP}({\mathcal{G}}_{{\mathcal{C}}})$, where ${\mathcal{C}}$ is a cycle of cliques. To this end, next we introduce a lifting operation for the multilinear polytope that is key to our characterization. For notational simplicity, for a node $v$, we use the notations $z_{v}$ (resp. $z_{v\cup e}$ for some $e\in{\mathcal{E}}$) and $z_{\{v\}}$ (resp. $z_{\{v\}\cup e}$ for some $e\in{\mathcal{E}}$), interchangeably.
Figure 1: Examples of cycle of cliques of length $m=3$ and $m=4$.
@@
Odd-cycle inequalities are a well-known class of valid inequalities for the Boolean quadric polytope [36]. These inequalities play an important role in characterizing the multilinear polytope of a cycle of cliques. We define these inequalities next; let $G=(V,E)$ be a graph. Padberg [36] introduced the Boolean quadric polytope of $G$, denoted by ${\rm BQP}(G)$, as follows:
@@
Let ${\mathcal{C}}=\{C_{1},\cdots,C_{m}\}$ be a cycle of cliques with $C_{i}\cap C_{i+1}=\{v_{i},\bar{v}\}$ for $i\in[m]$, where $C_{m+1}:=C_{1}$. Define the cycle $K:=\{\{v_{1},v_{i+1}\},\forall i\in[m]\}$, where $v_{m+1}:=v_{1}$. Denote by ${\mathcal{G}}_{C_{i}}$, $i\in[m]$ the complete hypergraph with node set $C_{i}$ and let ${\mathcal{G}}_{{\mathcal{C}}}=\cup_{i\in[m]}{{\mathcal{G}}_{C_{i}}}$. Then $\text{MP}({\mathcal{G}}_{{\mathcal{C}}})$ is obtained by juxtaposing inequalities defining $\text{MP}({\mathcal{G}}_{C_{i}})$ for $i\in[m]$, and the following inequalities:
@@
Define $C^{\prime}_{i}:=C_{i}\setminus\{\bar{v}\}$ for all $i\in[m]$. Denote by ${\mathcal{G}}_{C^{\prime}_{i}}$, the complete hypergraph with node set $C^{\prime}_{i}$ for $i\in[m]$ and define ${\mathcal{G}}_{{\mathcal{C}}^{\prime}}=\cup_{i\in[m]}{{\mathcal{G}}_{C^{\prime}_{i}}}$. By Proposition 5, to characterize $\text{MP}({\mathcal{G}}_{{\mathcal{C}}})$ it suffices to to characterize $\text{MP}({\mathcal{G}}_{{\mathcal{C}}^{\prime}})$. We next obtain the explicit description for $\text{MP}({\mathcal{G}}_{{\mathcal{C}}^{\prime}})$.
@@
The hypergraph ${\mathcal{G}}_{C^{\prime}_{1}}$ is the section hypergraph of ${\mathcal{G}}_{{\mathcal{C}}^{\prime}}$ induced by $C^{\prime}_{1}$. Denote by ${\mathcal{G}}_{2}$ the section hypergraph ${\mathcal{G}}_{{\mathcal{C}}^{\prime}}$ induced by $\cup_{i=2}^{m}{C^{\prime}_{i}}$. Notice that ${\mathcal{G}}_{2}=\cup_{i=2}^{m}{{\mathcal{G}}_{C^{\prime}_{i}}}\cup F_{m}$, where $F_{m}$ is the graph with node set $\{v_{1},v_{m}\}$ and edges set $\{\{v_{1},v_{m}\}\}$. We then have ${\mathcal{G}}_{{\mathcal{C}}^{\prime}}={\mathcal{G}}_{C^{\prime}_{1}}\cup{\mathcal{G}}_{2}$ and ${\mathcal{G}}_{C^{\prime}_{1}}\cap{\mathcal{G}}_{2}=\{v_{1},v_{m}\}$. Therefore, by Theorem 1 the multilinear set $S({\mathcal{G}}_{{\mathcal{C}}^{\prime}})$ decomposes into the multilinear sets $S({\mathcal{G}}_{C^{\prime}_{1}})$ and $S({\mathcal{G}}_{2})$. Next consider the hypergraph ${\mathcal{G}}_{2}$; denote by ${\mathcal{G}}_{3}$ the section hypergraph of ${\mathcal{G}}_{2}$ induced by $(\cup_{i=3}^{m}{C^{\prime}_{i}})\cup\{v_{1}\}$. Notice that ${\mathcal{G}}_{3}=\cup_{i=3}^{m}{{\mathcal{G}}_{C^{\prime}_{i}}}\cup F_{m}\cup F_{1}$, where $F_{1}$ is the graph with node set $\{v_{1},v_{2}\}$ and edges set $\{\{v_{1},v_{2}\}\}$. We then have ${\mathcal{G}}_{2}={\mathcal{G}}_{C^{\prime}_{2}}\cup{\mathcal{G}}_{3}$ and ${\mathcal{G}}_{C^{\prime}_{2}}\cap{\mathcal{G}}_{3}=\{v_{1},v_{2}\}$. Therefore, by Theorem 1, $S({\mathcal{G}}_{2})$ decomposes into $S({\mathcal{G}}_{C^{\prime}_{2}})$ and $S({\mathcal{G}}_{3})$. Applying this argument recursively, we conclude that $S({\mathcal{G}}_{{\mathcal{C}}^{\prime}})$ decomposes into $S({\mathcal{G}}_{C^{\prime}_{i}})$ for all $i\in[m]$ and $S(K)$, where the node set of $K$ is given by $\{v_{1},\cdots,v_{m}\}$ and the edge set of $K$ is given by $\{\{v_{i},v_{i+1}\},\forall i\in[m]\}$, where $v_{m+1}:=v_{1}$. First notice that the multilinear polytope $\text{MP}({\mathcal{G}}_{C^{\prime}_{i}})$, $i\in[m]$ is given by Proposition 1 as ${\mathcal{G}}_{C^{\prime}_{i}}$ is a complete hypergraph. Moreover, the multilinear polytope $\text{MP}(K)$ is given by Theorem 9 in [36] as $K$ is a chordless cycle; i.e., $\text{MP}(K)$ consists of odd-cycle inequalities (29) together with inequalities defining $\text{MP}(F_{i})$, $i\in[m]$ where the node set and the edge set of $F_{i}$ are given by $\{v_{i},v_{i+1}\}$ and $\{\{v_{i},v_{i+1}\}\}$, where $v_{m+1}:=v_{1}$. Since $F_{i}$ is a section hypergraph of ${\mathcal{C}}^{\prime}_{i}$ for all $i\in[m]$, we deduce that the inequalities defining $\text{MP}(F_{i})$ are implied by inequalities defining $\text{MP}({\mathcal{G}}_{C^{\prime}_{i}})$. Therefore, $\text{MP}({\mathcal{G}}_{{\mathcal{C}}^{\prime}})$ is defined by inequalities defining $\text{MP}({\mathcal{G}}_{C^{\prime}_{i}})$, $i\in[m]$ together with odd-cycle inequalities for $K$. Therefore, by Proposition 5, $\text{MP}({\mathcal{G}}_{{\mathcal{C}}})$ is defined by inequalities defining $\text{MP}({\mathcal{G}}_{C_{i}})$, $i\in[m]$ together with inequalities (6). ∎
@@
Henceforth, we refer to inequalities (6) as the lifted odd-cycle inequalities. We then define the multi-clique relaxation of the multilinear set, denoted by $\text{MP}^{\rm Mcl}({\mathcal{G}})$, as the polytope obtained by adding all lifted odd-cycle ineualities (6) corresponding to cycles of cliques of length at most $m\leq M$ to the clique relaxation $\text{MP}^{\rm cl}({\mathcal{G}})$. We now define our final LP relaxation which we refer to as the multi-clique LP:
@@
In [9, 17], the authors introduce odd $\beta$-cycle inequalities, a class of valid inequalities for the multilinear polytope. Let ${\mathcal{C}}$ be a cycle of cliques of length $m$ and let ${\mathcal{G}}_{{\mathcal{C}}}$ be the corresponding UGM hypergraph. Then it can be checked that there exist odd $\beta$-cycle inequalities of $\text{MP}({\mathcal{G}}_{{\mathcal{C}}})$ that are not implied by the clique relaxation $\text{MP}^{\rm cl}({\mathcal{G}}_{{\mathcal{C}}})$. However, by Proposition 6, thanks to inequalities (6), odd $\beta$-cycle inequalities are implied by the multi-clique relaxation $\text{MP}^{\rm mcl}({\mathcal{G}}_{{\mathcal{C}}})$. For a general UGM hypergraph ${\mathcal{G}}$, there may exist odd $\beta$-cycle inequalities are not implied by any multi-clique relaxation. However, as we demonstrate in Sections 3 an 4, for inference in binary UGMs, clique LP and multi-clique LP are often sharp and hence we do not explore other valid inequalities such as odd $\beta$-cycle inequalities.

### v2

@@ 2.4 The clique relaxation
Let ${\mathcal{G}}=\cup_{C\in{\mathcal{C}}}{{\mathcal{G}}_{C}}$ be a UGM hypergraph where ${\mathcal{C}}$ denotes the set of maximal clique of the binary UGM and ${\mathcal{G}}_{C}$ is a complete hypergraph with node set $C$. If ${\mathcal{C}}$ has the running intersection property, then $\text{MP}^{\rm cl}({\mathcal{G}})=\text{MP}({\mathcal{G}})$
@@
Denote by $C_{1},C_{2},\cdots,C_{m}$ a running intersection ordering of ${\mathcal{C}}$. Let ${\mathcal{G}}_{i}$ denote the complete hypergraph with node set $C_{i}$ for all $i\in\{1,\cdots,m\}$. Define ${\mathcal{G}}^{\prime}_{m}=\cup_{i=1}^{m-1}{{\mathcal{G}}_{i}}$. Then by definition of the running intersection ordering, ${\mathcal{G}}_{m}\cap{\mathcal{G}}^{\prime}_{m}$ is a section hypergraph of ${\mathcal{G}}_{i}$ for some $i\in\{1,\cdots,m-1\}$ and hence is a complete hypergraph. Hence, by Theorem 1 the set $S({\mathcal{G}})$ is decomposable into $S({\mathcal{G}}_{m})$ and $S({\mathcal{G}}^{\prime}_{m})$. By a recursive application of this argument $m$ times we conclude that $S({\mathcal{G}})$ is decomposable into multilinear sets $S({\mathcal{G}}_{i})$, $i\in\{1,\cdots,m\}$, implying $\text{MP}^{\rm cl}({\mathcal{G}})=\text{MP}({\mathcal{G}})$. ∎
@@ 2.5 The multi-clique relaxation
By Proposition 3, the clique relaxation is the strongest LP relaxation introduced so far. Yet by Proposition 4, this LP is guaranteed to solve the original problem if the set of cliques ${\mathcal{C}}$ has the running intersection property; a property that is often not present in applications. In this section we propose stronger LP relaxations for Problem B-UGM by constructing the multilinear polytope of a UGM hypergraph containing multiple cliques that do not have the running intersection property. More precisely, we consider a special structure, that we refer to as the lifted cycle of cliques, and obtain the multilinear polytope using disjunctive programming [2]. This structure appears in applications such as image restoration and decoding problems.
Consider the set of maximal cliques ${\mathcal{C}}:=\{C_{1},C_{2},\ldots,C_{m}\}$, where $m\geq 3$ and $|C_{i}|\geq 3$ for all $i\in[m]:=\{1,\cdots,m\}$. We say that ${\mathcal{C}}$ is a lifted cycle of cliques if $C_{i}\cap C_{i+1}=\{v_{i},\bar{v}\}$ for all $i\in[m]$, where $v_{i}\neq v_{j}$ for any $i\neq j\in[m]$ and where we define $C_{m+1}:=C_{1}$. It is simple to check if ${\mathcal{C}}$ is a lifted cycle of cliques, then it does not have the running intersection property. Figure 1 illustrates examples of cycles of cliques that appear in applications. The objective of this section is to characterize $\text{MP}({\mathcal{G}}_{{\mathcal{C}}})$, where ${\mathcal{C}}$ is a lifted cycle of cliques. To this end, next we introduce a lifting operation for the multilinear polytope which is the key to our characterization. For notational simplicity, for a node $v$, we use the notations $z_{v}$ (resp. $z_{v\cup e}$ for some $e\in{\mathcal{E}}$) and $z_{\{v\}}$ (resp. $z_{\{v\}\cup e}$ for some $e\in{\mathcal{E}}$), interchangeably.
Figure 1: Examples of lifted cycle of cliques of length $m=3$ and $m=4$.
@@
Odd-cycle inequalities are a well-known class of valid inequalities for the Boolean quadric polytope [38]. These inequalities play an important role in characterizing the multilinear polytope of a lifted cycle of cliques. We define these inequalities next; let $G=(V,E)$ be a graph. Padberg [38] introduced the Boolean quadric polytope of $G$, denoted by ${\rm BQP}(G)$, as follows:
@@
Let ${\mathcal{C}}=\{C_{1},\cdots,C_{m}\}$ be a lifted cycle of cliques with $C_{i}\cap C_{i+1}=\{v_{i},\bar{v}\}$ for $i\in[m]$, where $C_{m+1}:=C_{1}$. Define the cycle $K:=\{\{v_{1},v_{i+1}\},\forall i\in[m]\}$, where $v_{m+1}:=v_{1}$. Denote by ${\mathcal{G}}_{C_{i}}$, $i\in[m]$ the complete hypergraph with node set $C_{i}$ and let ${\mathcal{G}}_{{\mathcal{C}}}=\cup_{i\in[m]}{{\mathcal{G}}_{C_{i}}}$. Then $\text{MP}({\mathcal{G}}_{{\mathcal{C}}})$ is defined by the system comprising of inequalities defining $\text{MP}({\mathcal{G}}_{C_{i}})$ for $i\in[m]$, and the following inequalities:
@@
Define $C^{\prime}_{i}:=C_{i}\setminus\{\bar{v}\}$ for all $i\in[m]$. Denote by ${\mathcal{G}}_{C^{\prime}_{i}}$, the complete hypergraph with node set $C^{\prime}_{i}$ for $i\in[m]$ and define ${\mathcal{G}}_{{\mathcal{C}}^{\prime}}=\cup_{i\in[m]}{{\mathcal{G}}_{C^{\prime}_{i}}}$. By Proposition 5, to characterize $\text{MP}({\mathcal{G}}_{{\mathcal{C}}})$ it suffices to characterize $\text{MP}({\mathcal{G}}_{{\mathcal{C}}^{\prime}})$. We next obtain the explicit description for $\text{MP}({\mathcal{G}}_{{\mathcal{C}}^{\prime}})$.
@@
The hypergraph ${\mathcal{G}}_{C^{\prime}_{1}}$ is the section hypergraph of ${\mathcal{G}}_{{\mathcal{C}}^{\prime}}$ induced by $C^{\prime}_{1}$. Denote by ${\mathcal{G}}_{2}$ the section hypergraph ${\mathcal{G}}_{{\mathcal{C}}^{\prime}}$ induced by $\cup_{i=2}^{m}{C^{\prime}_{i}}$. Notice that ${\mathcal{G}}_{2}=\cup_{i=2}^{m}{{\mathcal{G}}_{C^{\prime}_{i}}}\cup F_{m}$, where $F_{m}$ is the graph with node set $\{v_{1},v_{m}\}$ and edge set $\{\{v_{1},v_{m}\}\}$. We then have ${\mathcal{G}}_{{\mathcal{C}}^{\prime}}={\mathcal{G}}_{C^{\prime}_{1}}\cup{\mathcal{G}}_{2}$ and ${\mathcal{G}}_{C^{\prime}_{1}}\cap{\mathcal{G}}_{2}=\{v_{1},v_{m}\}$. Therefore, by Theorem 1 the multilinear set $S({\mathcal{G}}_{{\mathcal{C}}^{\prime}})$ decomposes into the multilinear sets $S({\mathcal{G}}_{C^{\prime}_{1}})$ and $S({\mathcal{G}}_{2})$. Next consider the hypergraph ${\mathcal{G}}_{2}$; denote by ${\mathcal{G}}_{3}$ the section hypergraph of ${\mathcal{G}}_{2}$ induced by $(\cup_{i=3}^{m}{C^{\prime}_{i}})\cup\{v_{1}\}$. Notice that ${\mathcal{G}}_{3}=\cup_{i=3}^{m}{{\mathcal{G}}_{C^{\prime}_{i}}}\cup F_{m}\cup F_{1}$, where $F_{1}$ is the graph with node set $\{v_{1},v_{2}\}$ and edge set $\{\{v_{1},v_{2}\}\}$. We then have ${\mathcal{G}}_{2}={\mathcal{G}}_{C^{\prime}_{2}}\cup{\mathcal{G}}_{3}$ and ${\mathcal{G}}_{C^{\prime}_{2}}\cap{\mathcal{G}}_{3}=\{v_{1},v_{2}\}$. Therefore, by Theorem 1, $S({\mathcal{G}}_{2})$ decomposes into $S({\mathcal{G}}_{C^{\prime}_{2}})$ and $S({\mathcal{G}}_{3})$. Applying this argument recursively, we conclude that $S({\mathcal{G}}_{{\mathcal{C}}^{\prime}})$ decomposes into $S({\mathcal{G}}_{C^{\prime}_{i}})$ for all $i\in[m]$ and $S(K)$, where the node set of $K$ is given by $\{v_{1},\cdots,v_{m}\}$ and the edge set of $K$ is given by $\{\{v_{i},v_{i+1}\},\forall i\in[m]\}$, where $v_{m+1}:=v_{1}$. First notice that the multilinear polytope $\text{MP}({\mathcal{G}}_{C^{\prime}_{i}})$, $i\in[m]$ is given by Proposition 1 as ${\mathcal{G}}_{C^{\prime}_{i}}$ is a complete hypergraph. Moreover, the multilinear polytope $\text{MP}(K)$ is given by Theorem 9 in [38] as $K$ is a chordless cycle; i.e., $\text{MP}(K)$ consists of odd-cycle inequalities (28) together with inequalities defining $\text{MP}(F_{i})$, $i\in[m]$ where the node set and the edge set of $F_{i}$ are given by $\{v_{i},v_{i+1}\}$ and $\{\{v_{i},v_{i+1}\}\}$, where $v_{m+1}:=v_{1}$. Since $F_{i}$ is a section hypergraph of ${\mathcal{C}}^{\prime}_{i}$ for all $i\in[m]$, we deduce that the inequalities defining $\text{MP}(F_{i})$ are implied by inequalities defining $\text{MP}({\mathcal{G}}_{C^{\prime}_{i}})$. Therefore, $\text{MP}({\mathcal{G}}_{{\mathcal{C}}^{\prime}})$ is defined by inequalities defining $\text{MP}({\mathcal{G}}_{C^{\prime}_{i}})$, $i\in[m]$ together with odd-cycle inequalities for $K$. Therefore, by Proposition 5, $\text{MP}({\mathcal{G}}_{{\mathcal{C}}})$ is defined by inequalities defining $\text{MP}({\mathcal{G}}_{C_{i}})$, $i\in[m]$ together with inequalities (6). ∎
@@
Henceforth, we refer to inequalities (6) as the lifted odd-cycle inequalities. We then define the multi-clique relaxation of the multilinear set, denoted by $\text{MP}^{\rm Mcl}({\mathcal{G}})$, as the polytope obtained by adding all lifted odd-cycle inequalities (6) corresponding to cycles of cliques of length at most $m\leq M$ to the clique relaxation $\text{MP}^{\rm cl}({\mathcal{G}})$. We now define our final LP relaxation which we refer to as the multi-clique LP:
@@
In [9, 17], the authors introduce odd $\beta$-cycle inequalities, a class of valid inequalities for the multilinear polytope. Let ${\mathcal{C}}$ be a lifted cycle of cliques of length $m$ and let ${\mathcal{G}}_{{\mathcal{C}}}$ be the corresponding UGM hypergraph. Then it can be checked that there exist odd $\beta$-cycle inequalities of $\text{MP}({\mathcal{G}}_{{\mathcal{C}}})$ that are not implied by the clique relaxation $\text{MP}^{\rm cl}({\mathcal{G}}_{{\mathcal{C}}})$. However, by Proposition 6, thanks to inequalities (6), odd $\beta$-cycle inequalities are implied by the multi-clique relaxation $\text{MP}^{\rm mcl}({\mathcal{G}}_{{\mathcal{C}}})$. For a general UGM hypergraph ${\mathcal{G}}$, there may exist odd $\beta$-cycle inequalities are not implied by any multi-clique relaxation. However, as we demonstrate in Sections 3 an 4, for inference in binary UGMs, the clique LP and the multi-clique LP are often sharp and hence we do not explore other valid inequalities such as odd $\beta$-cycle inequalities.